OpenStudy (anonymous):

Which of the following is a factor of 2x3 – 2?

OpenStudy (anonymous):

\[ 2 x^3-2=2 (x-1) \left(x^2+x+1\right) \]

OpenStudy (anonymous):

Use \[ a^3- b^3 =(a-b)(a^2 + ab +b^2) \]

OpenStudy (anonymous):

Where are the following?